# A consensus-constrained parsimonious Gaussian mixture model for clustering hyperspectral images

### v1

@@ 3.2 Constrained parsimonious Gaussian mixture models
where $\hat{\cdot}$ denotes an initial or current parameter value, as relevant. In (3), the posterior probabilities of pixels in $B_{1}$ are considered with only one set of pixels $B_{2}$ as a negative constraint, however (3) can easily be extended to accommodate multiple such negative constraints. The remaining pixels in Figure 2 which are not in the highlighted blocks are considered as blocks with only one pixel and no negative constraints. The estimated model parameters $\hat{\bm{\mu}}_{g}$ and $\hat{\tau}_{g}$ are updated in the M-step of the first cycle as in PGMM(see McNicholas and Murphy, , 2008). Further details are in Appendix 9.
@@
In the E-step of the second cycle, the expected value of $\hat{{z}}_{B_{1},g}^{+\atop-}$ is computed as in (3), with $\hat{\bm{\mu}}_{g}$ and $\hat{\tau}_{g}$ updated to those from the M-step of the first cycle. The expected value of $\mathbf{u}_{n}\>,n=1,\ldots,N$, and the estimates of $\hat{\bm{\Lambda}}_{g}$ and $\hat{\bm{\Psi}}_{g}\>,g=1,\ldots,G$ are computed as in PGMM; details are reported in Appendix 9.
@@
The AECM algorithm moves between these cycles, iteratively updating the model parameters, until convergence. The values of $\hat{\mathbf{Z}}$ at convergence are the posterior probabilities of cluster membership for each pixel; labels can be obtained based, for example, on the maximum a posteriori values of $\hat{\mathbf{Z}}$. However, constrained-PGMM exhibits very long run times when fitted to the hyperspectral image data and could not be fitted to the entire $p=101$ variables of the motivating dataset in a computationally efficient manner.
@@ 3.3 Consensus-constrained parsimonious Gaussian mixture models
Unlike Russell et al., (2015), the proposed ccPGMM takes a divide-and-conquer approach. In ccPGMM, a constrained-PGMM is fitted to $M$ subsets (indexed by $m$) of $d<p$ randomly selected variables from the high dimensional NIR spectrum. The $M$ posterior probabilities which account for the uncertainty of the cluster memberships of the $N$ pixels are then consolidated to provide the final consensus clustering solution as follows. An $N\times N$ similarity matrix $\mathbf{S}^{m}$ for $m=1,\ldots,M$ is computed based on the estimated posterior probabilities $\hat{\mathbf{Z}}^{m}$ available on convergence of the AECM after fitting the constrained-PGMM to the subset $m$. Each entry of $\mathbf{S}^{m}$ is a similarity score $S_{ij}^{m}$ computed as
@@ 3.4 Technical details
In terms of specifying the constraints, there are technically no limits on the number of pixels that can be used. The size of the blocks of pixels (constraints) highlighted in Figure 2 can vary and the number of pixels in one block need not be the same as the number of pixels in the other blocks. Usually, the hyperspectral images of food samples are plotted as greyscale images, and the maximum region of pixels that are distinguishable as the background and as the relevant sample under study are selected as constraints. Increasing the number of positive constraints and the number of pixels within the positive constraints does not significantly increase the computational cost of fitting the ccPGMM, as in Melnykov et al., (2016). However, an increase in the number of negative constraints will result in more complex posterior probability calculations in (3), negatively impacting the computational cost. Investigations on the choice of constrained pixels are in sections 4 and 5.
@@ 4 Simulation studies
Simulation studies are conducted to assess the performance of ccPGMM and compare it to the performance of the threshold approach and existing state-of-the-art methods such as DBSCAN (Khan et al., , 2014; Hahsler et al., , 2019), Gaussian mixture models (GMM, Scrucca et al., , 2016; McNicholas, , 2016; Bouveyron et al., , 2019), PGMM (McNicholas and Murphy, , 2008) and consensus PGMM (cPGMM) i.e., ccPGMM with no constraints. The convergence of each model-based clustering approach is assessed with the default parameter settings detailed in the respective R packages. Four scenarios are considered: well-separated clusters (scenario 1), heavily overlapping clusters (scenario 2), mildly overlapping clusters (scenario 3), and synthetic puffed cereal data (scenario 4). For each scenario, five datasets are simulated.
@@
In scenarios 1–3, each simulated dataset contains three simulated hyperspectral images (one for each cereal type), giving a total of $N=7825$ pixels, where each pixel has an associated $p=101$ spectrum of strongly correlated variables. The three images are simulated according to labels resulting from applying the threshold approach to three of the hyperspectral images in the motivating dataset detailed in Section 2. The $p$ variables for each of the three types of cereal and the background pixels are generated from different factor analysis models with $q=3$ and parameters $\bm{\mu}_{g}$, $\bm{\Lambda}_{g}$, and $\bm{\Psi}_{g}$ for $g=1,\ldots,G=4$. To emulate the real data, a strong correlation is induced between the $p$ variables at each pixel by, for every set of $5$ consecutive variables, generating the associated values in the loadings matrix from a $\mathcal{N}_{q}(\mathcal{U}(0.3,0.9),0.03)$ and the values of $\mbox{diag}(\bm{\Psi}_{g})$ from a $\Gamma(\mathcal{U}(0,0.1),1)$. The values of the means $\bm{\mu}_{g}$ for $g=1,\ldots,G$ are used to control the degree of overlap of the clusters in each scenario; details are given in the respective subsections of Section 4. To determine the constraints, the greyscale images of the simulated hyperspectral images are examined to decide on the sets of pixels to be used as constraints, as shown in Figure 3. Here, this resulted in $2,956$ pixels of the $N=7825$ pixels $(37.7\%)$ being used to inform the constraints. The $1,356$ pixels in the blue blocks (background) must be clustered together, as must the $750$ pixels in the yellow block (wheat), the $450$ pixels in the green block (corn), and the $400$ pixels in the red block (rice). These are the positive constraints. The pixels in one coloured block must not be clustered together with the pixels in another coloured block; these are the negative constraints.
@@
Figure 3: Greyscale image of three simulated hyperspectral images for one of the five well-separated cluster datasets. The pixels in the blue blocks are background, the pixels in the yellow block are wheat, the pixels in the green block are corn and the pixels in the red block are rice.
In scenario 4, a synthetic cereal dataset is considered which mirrors the properties of the motivating hyperspectral images. Each dataset contains $L=9$ simulated hyperspectral images (three for each cereal type) with $N=28,039$ and where each pixel has an associated $p=101$ length spectrum. The $p$ variables for each cereal type and background pixels are generated from the relevant factor analyzer model, fitted to the real data; details are discussed in Section 4.4. In this scenario, $12,215$ pixels of the $N=28,039$ pixels $(43.5\%)$ are selected as constraints by plotting the greyscale images of the synthetic puffed cereals. Of the $12,215$ pixels selected as constraints, $5,900$ are background, $3,130$ are wheat, $1,535$ are corn, and $1,650$ are rice. Further, to assess the impact of the proportion of pixels used as constraints on the clustering performance, a smaller set of $6,951$ pixels ($24.7\%$) are selected as constraints. Of the $6,951$ pixels, $3,801$ are background, $1,540$ are wheat, $783$ are corn and $827$ are rice. The set of the $43.5\%$ of pixels and the $24.7\%$ of pixels used as constraints are illustrated in Figures 13 and 14 respectively in Appendix 10.
@@
4.1 Scenario 1: well-separated clusters
To generate the well-separated clusters, for $g=1,\ldots,4$, the cluster mean parameters are generated as $\bm{\mu}_{g}\sim\mathcal{N}_{p}(a,0.5)$, where $a\in\{0,5,10,-5\}$; the $p$ variables of the background, wheat, corn, and rice pixels in each dataset are generated from the four well-separated factor analyzer models, respectively. For DBSCAN, the minimum points required in the neighborhood for core points are set to $2p$, and for the GMM, $G=4$ is determined by the setting. For the PGMM approaches, $G=4$ and $q=1$, where $q$ is informed by fitting PCA to the entire dataset and monitoring the proportion of variance explained by the components. For consensus based approaches, $M=\{10,25,50,100\}$ randomly selected subsets of variables are considered with the number of variables per subset $d=\{10,20\}$. The positive and negative constraints for ccPGMM are selected as shown in Figure 3.
@@ 4.1 Scenario 1: well-separated clusters
Figure 4: For each of the five simulated datasets with well-separated clusters, the ARI between the known labels and the clustering solutions of DBSCAN, GMM, and PGMM fitted on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
@@
Figure 5: For each of the five simulated datasets with well-separated clusters, the time taken to fit DBSCAN, GMM, and PGMM on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
@@
Figure 4 shows the adjusted Rand index(ARI, Hubert and Arabie, , 1985) between the known labels and the clustering solutions for the five simulated datasets in scenario 1 under the DBSCAN, GMM, PGMM, cPGMM, and ccPGMM approaches. Figure 5 shows the time taken in minutes to fit the methods. All approaches demonstrate strong clustering performance. In terms of computational cost, DBSCAN clusters each dataset in less than a minute while GMM and PGMM take an average of $27$ minutes and $15$ minutes respectively. The cPGMM and ccPGMM fitted with different settings of $M$ and $d$ have lower computational costs than GMM and PGMM, but similar to DBSCAN, with no loss of accuracy.
4.2 Scenario 2: heavily overlapping clusters
To generate data with heavily overlapping clusters, the cluster mean parameters are generated as $\bm{\mu}_{g}\sim\mathcal{N}_{p}(a,1.5)$, where $a\in\{1,1.5,3,-1.5\}$ for $g=1,\ldots,4$ respectively. Figures 6 and 7 show the ARI between the known pixel labels and the clustering solutions of DBSCAN, GMM, PGMM, cPGMM, and ccPGMM and the time taken in minutes by each method, respectively. DBSCAN clusters the datasets with an average ARI of $0.34$ in less than a minute on average. The GMM struggles to uncover the clustering structure. The PGMM approach performs competitively, achieving an average ARI of $0.40$ in $5$ minutes on average. The ccPGMM approach performs well in this scenario: with $M\geq 25$ and $d=20$ an average ARI of $0.66$ is achieved, which outperforms PGMM fitted to all $p$ variables. In terms of clustering performance, ccPGMM performs better than cPGMM across all settings of $M$ and $d$, in comparable time.
Figure 6: For each of the five datasets with heavily overlapping clusters, the ARI of DBSCAN, GMM, and PGMM fitted on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) with different settings of $M$ and $d$.
Figure 7: For each of the five datasets with heavily overlapping clusters, the time taken to fit DBSCAN, GMM, and PGMM on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) with different settings of $M$ and $d$.
4.3 Scenario 3: mildly overlapping clusters
@@
The $p$ variables of the background, wheat, corn, and rice pixels in each dataset are generated from four mildly overlapping factor analyzer models, with the cluster mean parameters generated as $\bm{\mu}_{g}\sim\mathcal{N}(a,1.5)$ where $a=\{0,2,5,-2\}$ for $g=1,\ldots,4$, respectively. Figures 8 and 9 show the ARI of the cluster solutions of DBSCAN, GMM, PGMM, cPGMM, and ccPGMM fitted on the five datasets with mildly overlapping clusters and the time taken in minutes to fit them. While GMM again struggles to uncover the clustering structure, DBSCAN and PGMM fitted to all $p$ variables clustered the pixels with an average ARI of $0.42$ and $0.64$ respectively. In terms of cPGMM and ccPGMM, Figure 8 shows that increasing $d$ from $10$ to $20$ significantly improves the clustering performance of both methods, with the performance of ccPGMM being the best among the approaches considered. Run times are comparable overall, with cPGMM with the highest values of $M$ and $d$ intuitively taking the longest. The ccPGMM approach with $M\geq 10$ and $d=20$ is competitively computationally efficient with an average ARI $>0.80$, which has again a much better performance than PGMM fitted to all $p$ variables.
@@
Figure 8: For each of the five datasets with mildly overlapping clusters, the ARI of DBSCAN, GMM, and PGMM fitted on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
@@
Figure 9: For each of the five datasets with mildly overlapping clusters, the time taken to fit DBSCAN, GMM, and PGMM on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.

### v2

@@ 3.2 Constrained parsimonious Gaussian mixture models
In (4), only one negative constraint is considered, however (4) can easily be extended to accommodate multiple such negative constraints. The remaining pixels in Figure 2 which are not in the highlighted blocks have no constraints. To complete the first cycle of the AECM, the model parameters $\bm{\Theta}_{1}$ are updated in the conditional maximization (CM) step with $\bm{\Theta}_{2}$ held fixed as in PGMM (see McNicholas and Murphy, , 2008); further details are given in Appendix B.
@@
In the second cycle, both component indicators $\mathbf{Z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{n}\}$ and latent factors $\mathbf{U}=\{\mathbf{u}_{1},\ldots,\mathbf{u}_{n}\}$ are assumed to be the missing data. In the E-step of the second cycle, the expected value of $\hat{{z}}_{B_{1},g}^{+\atop-}$ is computed as in (4), with $\bm{\Theta}_{1}$ updated to the estimates from the CM-step of the first cycle. The expected value of $\mathbf{u}_{n}(n=1,\ldots,N$) and estimates of $\bm{\Theta}_{2}$ are computed as in PGMM; details are reported in Appendix B.
@@
The AECM algorithm moves between these two cycles, iteratively updating the model parameters, until convergence. The values of $\hat{\mathbf{Z}}$ at convergence are the posterior probabilities of cluster membership for each pixel; labels can be obtained based, for example, on the maximum a posteriori values of $\hat{\mathbf{Z}}$. However, constrained-PGMM exhibits very long run times when fitted to the hyperspectral image data and could not be fitted to the entire $p=101$ variables of the motivating dataset in a computationally efficient manner.
@@ 3.3 Consensus-constrained parsimonious Gaussian mixture models
Unlike Russell et al., (2015), the proposed ccPGMM takes a divide-and-conquer approach. In ccPGMM, a constrained-PGMM is fitted to $M$ subsets (indexed by $m$) of $d<p$ randomly selected variables from the high-dimensional NIR spectrum. The $M$ posterior probabilities which account for the uncertainty of the cluster memberships of the $N$ pixels are then consolidated to provide the final consensus clustering solution. An $N\times N$ similarity matrix $\mathbf{S}^{m}$ for $m=1,\ldots,M$ is computed based on the estimated posterior probabilities $\hat{\mathbf{Z}}^{m}$ available on convergence of the AECM after fitting the constrained-PGMM to the subset $m$. Each entry of $\mathbf{S}^{m}$ is a similarity score $S_{ij}^{m}$ computed as
@@ 3.4 Technical details
In terms of specifying the constraints, there are technically no limits on the number of pixels that can be used. The size of the blocks of pixels (constraints) highlighted in Figure 2 can vary and the number of pixels in one block need not be the same as the number of pixels in the other blocks. Usually, the hyperspectral images of food samples are plotted as greyscale images, and the maximum region of pixels that are distinguishable are selected as constraints. Increasing the number of positive constraints and the number of pixels within the positive constraints does not significantly increase the computational cost of fitting the ccPGMM, as in Melnykov et al., (2016). However, an increase in the number of negative constraints will result in more complex posterior probability calculations in (4), negatively impacting the computational cost. Investigations on the choice of constrained pixels are in Sections 4 and 5.
@@ 4 Simulation studies
Simulation studies are conducted to assess the performance of ccPGMM and compare it to the performance of the threshold approach and existing state-of-the-art methods such as DBSCAN (Khan et al., , 2014; Hahsler et al., , 2019), Gaussian mixture models (GMM, Scrucca et al., , 2016; McNicholas, , 2016; Bouveyron et al., , 2019), PGMM (McNicholas and Murphy, , 2008) and consensus PGMM (cPGMM) i.e., ccPGMM with no constraints. For DBSCAN (Hahsler et al., , 2019), the minimum points required in the neighborhood for core points was set to $2p$, and the size of the neighborhood was decided from the $k$-nearest neighbours distance plot (Peterson, , 2009). For the GMM, $G=4$ as determined by the setting and the full suite of models available in the mclust (Scrucca et al., , 2016) R package were fitted. For PGMM, the unconstrained model of McNicholas et al., (2023) was fitted with $G=4$ and $q$ set by considering the proportion of variance in the dataset explained by its principal components. Convergence in the GMM and PGMM approaches was assessed using the default settings detailed in their respective R packages. Consensus-based approaches were implemented with $G=4$ and $q$ set by monitoring the proportion of variance explained by the data principal components, considering $M=\{10,25,50,100\}$ randomly selected subsets with $d=\{10,20\}$ variables.
@@
Four scenarios are considered: low overlap clusters (scenario 1), mild overlap clusters (scenario 2), high overlap clusters (scenario 3), and synthetic puffed cereal data (scenario 4). For each scenario, five datasets are simulated. In scenarios 1–3, each simulated dataset contains three simulated hyperspectral images (one for each cereal type), for a total of $N=7825$ pixels, where each pixel has an associated spectrum of $p=101$ strongly correlated variables. The three images are simulated according to labels resulting from applying the threshold approach to three of the hyperspectral images in the motivating dataset detailed in Section 2. The $p$ variables for each of the three types of cereal and the background pixels are generated from different factor analysis models with $q=3$ and parameters $\bm{\mu}_{g}$, $\bm{\Lambda}_{g}$, and $\bm{\Psi}_{g}$ for $g=1,\ldots,G$ where in this setting $G=4$. To emulate the real data, a strong correlation is induced between the $p$ variables at each pixel by, for every set of $5$ consecutive variables, generating the associated values in the loadings matrix from a $\mathcal{N}_{q}(\mathcal{U}(0.3,0.9),0.03)$ and the values of $\mbox{diag}(\bm{\Psi}_{g})$ from a $\mathrm{Ga}(\mathcal{U}(0,0.1),1)$ where $\mathcal{U}$ and $\mathrm{Ga}$ denote the uniform and gamma distributions, respectively. The values of the means $\bm{\mu}_{g}$ for $g=1,\ldots,G$ are used to control the degree of overlap of the clusters in each scenario; details are given in the respective subsections of Section 4. To determine the constraints, the greyscale images of the simulated hyperspectral images are examined to decide on the sets of pixels to be used as constraints, as shown in Figure 3. Here, this resulted in $2,956$ pixels of the $N=7825$ pixels $(37.7\%)$ being used to inform the constraints. The $1,356$ pixels in the blue blocks must be clustered together, as must the $750$ pixels in the yellow block, the $450$ pixels in the green block, and the $400$ pixels in the red block. These are the positive constraints. The pixels in one coloured block must not be clustered together with the pixels in other coloured blocks; these are the negative constraints.
@@
Figure 3: Greyscale image of three simulated hyperspectral images for one of the five well-separated cluster datasets. The pixels in the blue blocks must be clustered together, as must the pixels in the yellow, green, and red blocks. The pixels in one coloured block must not be clustered with the pixels in other coloured blocks.
In scenario 4, a synthetic cereal dataset is considered which mirrors the properties of the motivating hyperspectral images. Each dataset contains $L=9$ simulated hyperspectral images (three for each cereal type) with $N=28,039$ and where each pixel has an associated $p=101$ length spectrum. The $p$ variables for each cereal type and background pixels are generated from the relevant factor analysis model, fitted to the real data; details are discussed in Section 4.4. In this scenario, $12,215$ pixels of the $N=28,039$ pixels $(43.5\%)$ are selected as constraints by plotting the greyscale images of the synthetic puffed cereals. Of the $12,215$ pixels selected as constraints, $5,900$ are in the blue blocks, $3,130$ are in the yellow blocks, $1,535$ are in the green blocks, and $1,650$ are in the red blocks. Further, to assess the impact of the proportion of pixels used as constraints on the clustering performance, a smaller set of $6,951$ pixels ($24.7\%$) are selected as constraints. Of the $6,951$ pixels, $3,801$ are in the blue blocks, $1,540$ is in the yellow blocks, $783$ are in the green blocks and $827$ are in the red blocks. The set of the $43.5\%$ of pixels and the $24.7\%$ of pixels used as constraints are illustrated in Figures 14 and 15 respectively in Appendix C.
@@
4.1 Scenario 1: low overlap clusters
To generate the low overlap clusters, the cluster mean parameters are generated as $\bm{\mu}_{g}\sim\mathcal{N}_{p}(a,0.5)$ for $g=1,\ldots,4$, where $a\in\{-5,0,5,10\}$. The $p$ variables of the background, wheat, corn, and rice pixels in each dataset are then generated from the four low overlapping factor analysis models.
@@ 4.1 Scenario 1: low overlap clusters
Figure 4: For each of the five simulated datasets with low overlap clusters, the ARI between the known labels and the clustering solutions of DBSCAN, GMM, and PGMM fitted on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
@@
Figure 5: For each of the five simulated datasets with low overlap clusters, the time taken to fit DBSCAN, GMM, and PGMM on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
@@
Figure 4 shows the adjusted Rand index (ARI, Hubert and Arabie, , 1985) between the known labels and the clustering solutions for the five simulated datasets in scenario 1 under the DBSCAN, GMM, PGMM, cPGMM, and ccPGMM approaches. For the parsimonious models, PCA suggested $q=1$. Figure 5 shows the time taken in minutes to fit the methods. All approaches demonstrate strong clustering performance. In terms of computational cost, DBSCAN clusters each dataset in less than a minute while GMM and PGMM take an average of $21$ minutes and $12$ minutes respectively. The cPGMM and ccPGMM fitted with different settings of $M$ and $d$ have lower computational costs than GMM and PGMM, but similar to DBSCAN, with no loss of accuracy.
4.2 Scenario 2: mild overlap clusters
@@
The $p$ variables of the background, wheat, corn and rice pixels in each dataset are generated from four mildly overlapping factor analysis models, with the cluster mean parameters generated as $\bm{\mu}_{g}\sim\mathcal{N}_{p}(a,1)$ where $a=\{-2.5,0,2.5,5\}$ for $g=1,\ldots,4$. Figures 6 and 7 show the ARI of the cluster solutions of DBSCAN, GMM, PGMM, cPGMM, and ccPGMM fitted on the five datasets with mild overlap clusters and the time taken in minutes to fit them, respectively. For the parsimonious models, PCA suggested $q=1$. While GMM struggles to uncover the clustering structure, DBSCAN, and PGMM fitted to all $p$ variables clustered the pixels with an average ARI of $0.48$ and $0.99$, respectively. In terms of cPGMM and ccPGMM, Figure 6 shows that increasing $d$ from $10$ to $20$ improves the clustering performance of both methods, with the performance of ccPGMM being preferable. Run times are comparable overall, with cPGMM with the highest values of $M$ and $d$ intuitively taking the longest. The ccPGMM approach with $M\geq 10$ and $d=20$ is competitively computationally efficient with an average ARI $>0.97$, a similar performance to PGMM fitted to all $p$ variables, but at a reduced computational cost.
@@
Figure 6: For each of the five datasets with mild overlap clusters, the ARI of DBSCAN, GMM, and PGMM fitted on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
@@
Figure 7: For each of the five datasets with mild overlap clusters, the time taken to fit DBSCAN, GMM, and PGMM on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) fitted with different settings of $M$ and $d$.
4.3 Scenario 3: high overlap clusters
To generate data with high overlap clusters, the cluster mean parameters are generated as $\bm{\mu}_{g}\sim\mathcal{N}_{p}(a,1.5)$, where $a\in\{-1.25,0,1.25,2.5\}$ for $g=1,\ldots,4$ respectively. Figures 8 and 9 show the ARI between the known pixel labels and the clustering solutions of DBSCAN, GMM, PGMM, cPGMM, and ccPGMM and the time taken in minutes by each method, respectively. For the parsimonious models, PCA suggested $q=1$. DBSCAN clusters the datasets with an average ARI of $0.38$ in less than a minute on average. The GMM struggles to uncover the clustering structure. The PGMM approach performs competitively, achieving an average ARI of $0.42$ in $6$ minutes on average. The ccPGMM approach performs well in this scenario: with $M\geq 25$ and $d=20$ an average ARI of $0.64$ is achieved, which outperforms PGMM fitted to all $p$ variables. In terms of clustering performance, ccPGMM performs better than cPGMM across all settings of $M$ and $d$, in comparable time.
Figure 8: For each of the five datasets with high overlap clusters, the ARI of DBSCAN, GMM, and PGMM fitted on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) with different settings of $M$ and $d$.
Figure 9: For each of the five datasets with high overlap clusters, the time taken to fit DBSCAN, GMM, and PGMM on $p=101$ variables (first three panels from the left) and cPGMM and ccPGMM (last two panels on the right) with different settings of $M$ and $d$.